OpenStudy (anonymous):

need help with ONE algebra question, will give best response if you answer correctly

Directrix (directrix):

2x + 3y = 6 2x - 3y = 6 --> multiply each member of the equation by (1) ------------

Directrix (directrix):

to get: 2x + 3y = 6 -2x + 3y = - 6 =========== Add 0 + 6 y = 0 ---->@iappreciateyourhelp Do you agree so far? Questions?

Directrix (directrix):

@iappreciateyourhelp

Directrix (directrix):

Then, 0 + 6 y = 0 6y = 0 Divide both sides by 6 to get: y = 0

OpenStudy (anonymous):

I got it now, thank you!

Directrix (directrix):

You are welcome. Please ask any questions you have.

OpenStudy (anonymous):

gave you best response for being so helpful :)

OpenStudy (anonymous):

that's all the questions I have for now but I will become a fan of yours if I have any future ones.

Directrix (directrix):

Thank you. What I had intended to convey is to ask if you had questions about this particular problem.

OpenStudy (anonymous):

oh no I'm okay now, you've done very well explaining and going through it so now I understand it better